### v4

# Era Splitting - Invariant Learning for Decision Trees

Abstract
Real-life machine learning problems exhibit distributional shifts in the data from one time to another or from one place to another. This behavior is beyond the scope of the traditional empirical risk minimization paradigm, which assumes i.i.d. distribution of data over time and across locations. The emerging field of out-of-distribution (OOD) generalization addresses this reality with new theory and algorithms which incorporate environmental, or era-wise information into the algorithms. So far, most research has been focused on linear models and/or neural networks. In this research we develop two new splitting criteria for decision trees, which allow us to apply ideas from OOD generalization research to decision tree models, namely, gradient boosting decision trees (GBDT). The new splitting criteria use era-wise information associated with the data to grow tree-based models that are optimal across all disjoint eras in the data, instead of optimal over the entire data set pooled together, which is the default setting. In this paper, two new splitting criteria are defined and analyzed theoretically. Effectiveness is tested on four experiments, ranging from simple, synthetic to complex, real-world applications. In particular we cast the OOD domain-adaptation problem in the context of financial markets, where the new models out-perform state-of-the-art GBDT models on the Numerai data set. The new criteria are incorporated into the Scikit-Learn code base and made freely available online.
1 Introduction
Gradient boosting regression trees ([30]) and descendent algorithms such as those implemented via the Light Gradient Boosting Machine ([32]), XGBoost ([26]), and Scikit-learn’s HistGradientBoostingRegressor ([38]) are considered state of the art for many real-world supervised learning problems characterized by medium data size, tabular format, and/or low signal-to-noise ratio ([31, 28]). One such data science problem occurs with predicting future returns in the capital markets. When this problem is framed as a standard supervised learning problem, the target variable is assumed to be a continuous score, representing the expected return (alpha) for a particular stock. The input data is the available information about that stock at that particular time. Each stock’s input data and associated target for each date are represented as one row in a data set. This problem is generally referred to as predicting the cross-sectional returns of the equity market.
Supervised learning models in their original form are not aware of the date information presented in the data. One row of input data is associated with one target. Samples are assumed to be drawn I.I.D. from the training and test sets, per the empirical risk minimization (ERM) principal ([40]). While in reality $N$ samples are drawn simultaneously from the data set, where $N$ is the number of stocks in the tradable universe. The predictions for all $N$ stocks are created together at the same time. These $N$ data points may not be completely I.I.D. either, as there are measurable positive correlations among stocks in the stock market over any particular time period. Indeed these ideas are the foundation of the capital asset pricing model (CAPM) ([27]) and modern portfolio theory ([35]), as well as the principles of risk factor models such as the Barra risk model ([24]). Thus, the cross-sectional stock prediction problem violates many of the assumptions of the ERM pricniple that determine the efficacy of the model predictions.
The discrepancy between the foundations of the ERM principle and the realities of ML application fits into the contemporary literature in the academic field of out-of-distribution generalization (OOD) research ([34]). This field of study recognizes that the ERM principle does not hold true for many, if not all, real-life applications. It does not assume that the data distribution of our training and test sets are identical. Instead, the framework assumes a certain distributional shift among disjoint data sets comprising the training data. Models are designed to account for these shifts. In the literature, this concept is often described in terms of environments or domains. In the context of this research, an environment is a domain, is an era, and is a date/location. Drawing data from different environments is analogous to drawing data from different dates or locations. Only data drawn from the same environment follows the same distribution, but across environments the distribution changes. Data sets are composed of data from several environments.
Environments are simply subsets of the entire data set. There is some distributional shift in the data-generating process from one environment to another. The purpose of this is to recognize that there may be spurious signals that work in a single or even a group of environments. It is even shown in ([37]) that these spurious signals can exist when many environments are pooled together. But these spurious signals change or disappear from one environment to another. These spurious signals usually will present themselves as simpler interpretations of the problem, so a naive model will quickly latch onto them. Think of the cow on grass and camel on sand image recognition problem ([36]). During training an image classification task for predicting the label of the animal in an image, cows are always displayed on green grass backgrounds and camels always appear on sandy backgrounds. Naive ML models will simply learn to associate the color green with a cow and the color yellow with a camel. Of course, this is a severe over-simplification of the problem. During evaluation, a cow appearing on a sandy background is classified as a camel, and a camel on a grassy background is classified as a cow. A model which latches onto these spurious signals will perform well in-sample, but have severe flaws when evaluated out-of-sample (OOS).
Fortunately there also exist invariant signals which work in all environments. These signals can often be much more complex than the spurious signals, so naive models are not likely to learn these invariant signals as readily. In the above example, the invariant signal is the cow or camel that is actually present in the image. If an ML model truly understands what a camel or cow looks like, then it will have no problem predicting correctly no matter the background.
Figure 1: Financial data experience distribution shifts over time due to changing macro-economic variables and current events. OOD algorithms endeavor to uncover signals which are present in all eras (environments) instead of spurious signals only present in some.
The purpose of this field of research is to design ML procedures which allow models to ignore the spurious and learn the invariant signals in data. In the work of ([21]), which is considered foundational in the field, the authors define their problem as a constrained optimization problem and convert it into a loss function that can be used in gradient descent-style training procedures. The loss function combines that tradition ERM risk term, which minimizes the average error over all environments, and an additional penalty on the magnitude of the gradients of the error in each environment. The gradient norm term is used to measure the optimality of the solution in each environment. Considerable theoretical and data analysis is performed around this idea, and the technique is implemented via neural networks and shows that it works on synthetic data problems designed for this purpose, such as the Colored MNIST problem.
A similarly influential work ([37]) also analyzes the phenomenon of spurious and invariant signals existing simultaneously in data, and how to disentangle the signals. Here, the authors look at the loss surface, and describe the co-existence of local minima on the surface where some minima correspond to spurious predictors and others correspond to the invariant predictors. The insight here is to realize that spurious minima occur is some environments and not others, but invariant minima occur in all environments in the same location. The authors extend this reasoning to the directions of the gradients of the loss function with respect to the model parameters. The directions of the gradients for invariant minima should agree across all environments but the directions for spurious signals will disagree between environments. In order to realize this understanding practically, there is an alteration made to the traditional gradient descent algorithm which is called the AND-mask. Before propagating the model weights down the gradients at each iteration, first the gradients are computed using data from each environment separately. Gradient descent is only propagated in the directions of the gradients which agree over all the environments, and not along those that disagree. There is also a variable to adjust the level of agreement required. With this technique the authors show how neural network models can effectively extract invariant predictors from data while ignoring spurious ones. It is important to note that the theoretical insights form the previous two references are limited to the linear settings, while the ML models function in high dimensions and non-linear signals.
Figure 2: A schematic example of splitting data at each tree node. The target values are stored inside each data point. The original setting pools all the data together. Era splitting computes split scores on a per era basis. The value of the child nodes, the directions of the splits and the original and era split criteria scores are indicated. Notice era splitting does not choose the same split as the original, since it wouldn’t improve impurity in any era.
The current research article investigates the OOD generalization problem in the context of gradient boosted decision trees (GBDTs), and related decision tree algorithms. In particular we focus on regression trees, which take a continuous variable as target, as opposed to trees for classification which take a class label as target. GBDTs implement decision trees as the base learner $f$. Trees are grown in a top-down approach, which is a cascading, step-wise process which intends to minimize the following objective function at each step, $t\in\{1,2,...\}$ ([25, 39, 30, 26]). The key piece of the objective function is the convex differentiable loss function $l(y_{i},\hat{y_{i}})$ which measures the loss between target $y_{i}$ and prediction $\hat{y}_{i}$. At each step $t$, the loss is computed between the true label and a correction $f_{t}(x_{i})$ to the prediction from the previous step. In the following equations $n$ is the total number of data points for input/target pairs $(x_{i},y_{i}),\forall i\in\{1,...,n\}$ and $\Omega$ is the regularizing function.
$\mathcal{L}^{(t)}=\sum_{i=1}^{n}l(y_{i},\hat{y}^{(t-1)}+f_{t}(x_{i}))+\Omega(f_{t})$ (1)
In practice the second order approximation to this objective is used in optimization, which puts the gradient in gradient boosting. In equation 2 $g_{i}=\partial_{y^{(t-1)}}l(y_{i},\hat{y}^{(t-1)})$ and $h_{i}=\partial^{2}_{y^{(t-1)}}l(y_{i},\hat{y}^{(t-1)})$. This derivation follows closely from ([26]) with more foundation in ([30]) and ([29]).
$\mathcal{L}^{(t)}=\sum_{i=1}^{n}l(y_{i},\hat{y}^{(t-1)})+g_{i}f_{t}(x_{i})+\frac{1}{2}h_{i}f_{t}^{2}(x_{i}))+\Omega(f_{t})$ (2)
Since the base learner $f$ is a decision tree, optimizing the decision tree function is equivalent to finding the optimal split point of the input data $x_{i}$ at each step. Each split point is defined by a single feature and a single value of that feature’s data. Starting at the the root node, which contains all the training data, the optimal split point is chosen. All the data with that feature greater than or equal to the split value gets partitioned into the left child node, and the data points with the feature less than the split value go to the right child node. This is a decision tree of depth 1. In successive steps the child nodes become parent nodes, and the data in those nodes is split in the same manner as the root node. The function that determines the optimal split point at each node is called the splitting criterion, which measures the reduction in impurity (also called information gain) induced by the split. The splitting criterion assigns a score to each possible split of the data at each node. The split is chosen which has the highest score, as this is the optimal split, resulting in the purest child nodes. Impurity is measured in the targets, or labels. Good splits create child nodes with the low impurity. For an introduction to decision trees and a review of the popular splitting criteria and impurity measures, refer to the text ([39]). In GBDTs for regression, the popular spitting criterion implements a custom measure of impurity described in ([26]), $\mathcal{L}_{\text{imp}}$, in equation 3.
$\mathcal{L}_{\text{imp}}=-\frac{1}{2}\sum_{j=1}^{T}\frac{\left(\sum_{i\in I_{j}}g_{i}\right)^{2}}{\sum_{i\in I_{j}}h_{i}+\lambda}-\gamma T$ (3)
$\mathcal{L}_{\text{imp}}$ computes the quality of an entire tree structure, where $g_{i}$ and $h_{i}$ are the gradients and hessians in leaf node $j\in\{1,...,T\}$ of the convex differentiable loss function. It is analogous to computing the variance in the node, but this function attains a maximum for minimum variance. In practice it isn’t feasible to evaluate every possible tree structure, but instead the iterative process grows a tree from the root node by selecting the best split based on the decrease in impurity from the parent to the child nodes.
$\mathcal{L}_{\text{split}}=\frac{1}{2}\left[\frac{\left(\sum_{i\in I_{L}}g_{i}\right)^{2}}{\sum_{i\in I_{L}}h_{i}+\lambda}+\frac{\left(\sum_{i\in I_{R}}g_{i}\right)^{2}}{\sum_{i\in I_{R}}h_{i}+\lambda}-\frac{\left(\sum_{i\in I}g_{i}\right)^{2}}{\sum_{i\in I}h_{i}+\lambda}\right]-\gamma$ (4)
Equation 4 is the original splitting criterion. $I$, $I_{L}$, and $I_{R}$ are data set identifiers corresponding to the data of the parent node, the left child node, and the right child node respectively. While equation 3 sums over all $T$ leaf nodes, the splitting criterion only evaluates the impurity of the current child nodes compared to the parent node. For GBDTs, implementing regression with the mean-squared error loss, a common simplification is that the Hessians are constant and equal to 1. Highest scores imply the greatest reduction in variance in the child nodes.
In the basic setting, GBDTs pool all the training data together when computing the splitting criterion. If the training data inherently comes from separate environments (eras), as defined in the OOD literature, what is lost by pooling all the training data together to find the best split? Would it be better to employ a splitting criterion that treats data from each environment (era) separately? This research endeavors to show that the answer to this question is, ”yes”.
A new splitting criterion called era splitting is proposed in section 2. It combines the impurity reduction calculation from each era separately into one final new criterion, which is a smooth average of the impurity reduction of the era-wise splits. It captures the essence of invariance from the IRM paradigm applied to tree splitting criteria. An additional era-wise splitting criterion is developed in section 3, which is designed after the AND-mask, and looks for splits which maximize the agreement in the direction of the predictions implied by the split for each era (environment). In section 4, the theoretical interpretation is expounded briefly, giving three new propositions. Section 5 then describes the experimental design used to validate our new splitting criteria. The experiments include a one-dimesnional toy model aimed at verifying the objectives of our new splitting, a high-dimensional synthetic dataset proposed for the same purpose in ([37]), and two real-world empirical studies. The first is the Camelyon17 dataset ([23]) breast cancer detection domain generalization problem, the second is financial stock market cross-sectional return problem provided by the Numerai data science tournament. In section 6, the results from the experiments are presented. Finally section 7 provides interpretation of the results and discussion about future work.
2 Era Splitting
The term era is used in the same sense as the term environment is used in the references ([21, 37]) and sometimes the term domain as in ([33]). An era of data is a disjoint set by which our training and test data sets are comprised of many eras of data. See definition 4.1 below. The general understanding is that each era contains data that occurs at certain distinct times and/or physical locations. In the context of the splitting criterion (4), the data set definitions $I$, $I_{L}$, and $I_{R}$ need to be refined when dealing with environments. With era splitting, each training data point comes from one of $M$ distinct eras. Thus there exists data sets $I^{j}$, $I_{L}^{j}$, and $I_{R}^{j}$, referring to data coming from the $j^{th}$ era, such that $I^{j}\in I$, $I_{L}^{j}\in I_{L}$, and $I_{R}^{j}\in I_{R}$ for $j\in\{1,2,...,M\}$. Using this convention, a new era-wise splitting criterion, which computes the information gain of a split for the $j^{th}$ era of data is given.
$\mathcal{L}_{\text{split}}^{j}=\frac{1}{2}\left[\frac{\left(\sum_{i\in I_{L}^{j}}g_{i}\right)^{2}}{\sum_{i\in I_{L}^{j}}h_{i}+\lambda}+\frac{\left(\sum_{i\in I_{R}^{j}}g_{i}\right)^{2}}{\sum_{i\in I_{R}^{j}}h_{i}+\lambda}-\frac{\left(\sum_{i\in I^{j}}g_{i}\right)^{2}}{\sum_{i\in I^{j}}h_{i}+\lambda}\right]-\gamma$ (5)
The era splitting criterion in its basic form is written as a mean average of the era-wise splitting criterion for each era of data.
$\mathcal{L}_{\text{era split}}=\frac{1}{M}\sum_{j=1}^{M}\mathcal{L}_{\text{split}}^{j}$ (6)
Figure 3: Degenerate split decisions induced by the traditional splitting criterion (Eq. 4) when viewed from an OOD setting. Each plot displays example data with 2 features from two eras (environments). The y-axis plots the gradients corresponding to each data point. On the left, the original splitting criterion chooses a split which doesn’t improve impurity in any era. The era splitting criterion (Eq. 8) chooses a split that improves impurity in both eras. On the right the original criterion chooses a split which results in conflicting directions, while directional era splitting (Eq. 11) chooses a split resulting in consistent directions in each era.
The era splitting criterion will have the highest score for the split which provides the greatest average decrease in impurity over all of the eras of data. Let us notice that this new criterion recovers equation 4 in the case where there is only one era of data. Therefore the era splitting criterion can be viewed as a generalization of the original splitting criterion. This criterion will be less-likely to favor splits which work really well in some eras but not in others. Splits are not considered which are undefined in any era. Potential split points must have data to the left and right of them to be considered as optimal. The reason for this choice is illustrated in figure 3. The plot on the left shows a data configuration where the original splitting criterion will choose a split which doesn’t induce any impurity decrease in any era of data. On the other era hand, era splitting chooses a split reducing impurity in both eras simultaneously.
A further generalization is made by way of a smooth max function, called the Boltzmann operator ([22]). The Boltzmann operator operates on an array of real numbers, and has one parameter, called $\alpha$. When $\alpha=0$, the Boltzmann operator recovers the arithmetic mean, when $\alpha=-\infty$ it recovers the $\min$, and when $\alpha=\infty$ it recovers the $\max$. To define the Boltzmann operator ($\mathcal{B}_{\alpha}$), consider $n$ real numbers, $x_{i}$ for $i\in\{1,...,n\}$.
$\mathcal{B}_{\alpha}(x_{1},...,x_{n})=\frac{\sum_{i=1}^{n}x_{i}\exp{(\alpha x_{i})}}{\sum_{i=1}^{n}\exp{(\alpha x_{i})}}$ (7)
The final version of the era splitting criterion is defined in equation 8, which is understood as the smooth maximum (minimum) of the era-wise impurity reduction over all of the eras, and is a generalization of the original splitting criterion in Eq 4.
$\mathcal{L}_{\text{era split}}^{\alpha}=\mathcal{B}_{\alpha}\left(\mathcal{L}_{\text{split}}^{1},...,\mathcal{L}_{\text{split}}^{M}\right)$ (8)
The default value of $\alpha$ is set to zero, which recovers equation 6. Conceptually, varying $\alpha$ toward $\infty$- ∞ will lead the criterion to prefer splits which increase the minimum information gain over all of the eras. This setting prefers splits that work in all eras. Conversely, varying $\alpha$ toward positive $\infty$ will cause the criterion to favor splits that improve the best single-era performance. In the prior case splits are preferred that have a more uniform affect on impurity decrease over all the eras. In the later case splits are preferred that improve our best performance in any era. In the context of learning an invariant predictor, negative values for $\alpha$ are preferred.
3 Directional Era Splitting
There is a subtle ambiguity that arises with equation 8, which requires the introduction of the concept of the direction implied by a split. A split implies a direction via the values of the child nodes. See figure 2. If the value of the left child node is larger than the value of the right child node, then the direction of this split is too the left. If the value of the right child node is larger, then the direction is to the right. Let the value of the left child node be $v_{l}$, and the value of the right child node be $v_{r}$, then without loss of generality, the direction of the split, $d_{\text{split}}$, is defined as the sign of the difference between the left child node value and the right child node value.
$d_{\text{split}}=\text{sign}(v_{l}-v_{r})$ (9)
In the traditional setting where all the training data is pooled together, there can be only one direction implied over the entire data set. The traditional splitting criterion (eq. 4) does not consider the direction of the split, since it is superfluous. It doesn’t matter whether the direction implied by the split is to the left or to the right, since there is only one era, just as long as the impurity is reduced. In era splitting, the split criterion is computed for each era separately, and then summarized with (the Boltzmann operator). A problem that arises here is that there could be splits that imply conflicting directions from one era to the next. We can describe the direction of a split in era $j$ by adapting equation 9 with era-wise descriptors, as follows.
$d_{\text{era split}}^{j}=\text{sign}(v_{l}^{j}-v_{r}^{j})$ (10)
The directions implied by the split for each era of data results in an array of length $M$, $\{d_{\text{era split}}^{0},d_{\text{era split}}^{1},...,d_{\text{era split}}^{M}\}$. Directional era splitting, $\mathcal{D}_{\text{era split}}$, prefers splits which have the most agreement in the direction across all eras.
$\mathcal{D}_{\text{era split}}=\frac{1}{M}\left|\sum_{j=0}^{M}d_{\text{era split}}^{j}\right|$ (11)
Each $d_{\text{era split}}^{j}$ will take a value of either $1$ or $-1$, depending on the direction of the split for that era. The sum is bounded on the interval $[-M,M]$. By taking the absolute value and dividing my $M$, $\mathcal{D}_{\text{era split}}$ is then bounded on the interval $[0,1]$. At the low end, with a value of 00, half of the directions go in one way while half go in the other way, the highest level of disagreement. At the high end, with a value of $1$, all the directions go the same way, the most agreement. When choosing the best split during tree growth, higher values of $\mathcal{D}_{\text{era split}}$ are preferred.
4 Theoretical Breakdown
This section expounds three propositions which are illustrative of the motivations behind era splitting. Here the focus is on era splitting (eq. 8), and not directional era splitting (eq. 11). The former being more anchored in traditional theory, and connected with the original criterion. The later is new and experimental, with impurity tied to the direction of the predictions in each era. Nevertheless, the following proposition 4.3 applies to both.
The problem with traditional splitting methods in the OOD setting is that the ”optimal” split can be degenerate in any or all of the eras (environments). The term, degenerate, indicates that the split doesn’t result in improved impurity in the child nodes compared to the parent for any particular era of the data set. This is a split that never would have been if the tree was grown on that era of data independently. Eras ( or environments ) usually occur independently in nature, such as disjoint time frames or locations. The left side of figure 3 shows a simple example of a degenerate split with a toy data set of 4 data points. The original split criterion favors the split (red dashed line) which effectively separates era 1 data from era 2 data. In each era, this split does not induce any reduction in impurity. The era splitting criterion, on the other hand, favors a split (green dotted line) which reduces impurity in both eras simultaneously.
Definition 4.1 (Era (Environment)).
Given the input-output training data $(X,Y)$ consisting of $N$ data points, an era (environment) of the data is a subset of the entire training set, $(X,Y)_{j}$ for $j\in\{1,..,M\}$, for which there are $M$ eras such that $\bigcup_{j}^{M}(X,Y)_{j}=(X,Y)$, and the eras do not intersect.
Definition 4.2 (Degenerate Splits).
A degenerate split is one which does not reduce impurity in any era of a training data.
Proposition 4.1.
The original splitting criterion defined in equation 4 can result in degenerate splits.
Proof.
Consider a data set consisting of two eras of data, two features and two data points, as defined in table 1, and appearing on the left plot of figure 3.
Feature 1 1, 2, 3, 4
Feature 2 1, 3, 2, 4
Era Identifier 0, 0, 1, 1
Target -1, -2, -3, -4
Prediction 0, 0, 0, 0
Gradient 1, 2, 3, 4
Table 1: Data for proposition 4.1.
The optimal split score for the original split criterion defined in equation 4 is equal to 2 and splits feature 1 between values 2 and 3. When computed independently over each era, this split results in an undefined score in both era 1 and era 2, since it does not split the data in either era. This split is degenerate.
This is one of the main conceptual motivations for the current research: it is possible that the original method misses, ignores, or misinterprets key environmental information. Era splitting addresses this problem in a very direct way: evaluate each split based on each environment separately. There is one configuration of era splitting which ensures the split points will always reduce impurity in every environment (era) of data simultaneously, never producing degenerate splits. This result is presented in proposition 4.2.
Proposition 4.2.
The era splitting criterion as defined in equation 8 with Boltzmann alpha parameter set to $\infty$- ∞ will never induce a degenerate split.
Proof.
The Boltzmann operator $\mathcal{B}_{\alpha}$ approaches the minimum operator as the value of $\alpha$ approaches infinity.
$\mathcal{B}_{\alpha}\to\min\ \text{as}\ \alpha\to-\infty$
In the context of era splitting (equation 8), the input to the operator are the split criterion scores from equation 4 computed in each era independently. The minimum is the lowest score from any era. In order to induce a split, this score must be greater than zero, as per the rules of node splitting. The worst improvement will still be positive. Thus, the era splitting criterion with Boltzmann alpha parameter of $\infty$- ∞ will only induce splits which reduce impurity in every era simultaneously.
A simple but important upshot of era splitting is that it always produces less impure splits, when viewed from the traditional setting. If we evaluate all splits using the original splitting criterion, other splits induced by other splitting criteria will, by definition, be worse (result in more impurity in the child nodes when all the data is pooled together as in the traditional setting). This means that trees grown with other splitting criteria will fit the training data to a lesser degree than the original. Impurity will not have been reduced as much. In-sample performance will be poorer. The hope is that this will result in better generalization to OOS performance. This is exactly the concept of regularization in ML. The experiments go on to confirm that this is exactly what is happening.
Proposition 4.3.
(Era Splitting as a Regularizer)
At each node, the era splitting (eq. 8 ) and directional era splitting (eq. 11) criteria will always choose splits with a score less than or equal to (worse than) the original criterion (eq. 4), when measured with the original criterion.
Proof.
Assume the original criterion (eq. 4) is used to find the optimal split point at a node. This split attains the highest score by definition. It is impossible for the era splitting (or any other) criterion to induce a split with a better score than the optimal split point, when measured using the original criterion. ∎
5 Experimental Methods
Four experiments have been designed to empirically validate the newly proposed splitting criteria. The first experiment is an original design, called the shifted sine wave, validating the era splitting models in a simple one-dimensional setting. The second and third experiments are binary classification problems from the OOD literature: the synthetic memorization data set presented in ([37]) and the Camelyon17 (CAncer MEtastases in LYmph nOdes challeNge) data set from ([23]). These two experiments confirm the new splitting criteria are effective in known OOD settings. The forth is the target application of this research, the Numerai data set.
The experiments follow a similar procedure for each. Random configurations are drawn from a grid of parameter values. The parameters include the number of boosting iterations, maximum number of leaves, learning rate, etc. All relevant parameters in Scikit-Learn’s HistGradientBoostingRegressor ([38]) (the baseline model) are available. The number of random configurations (between 15 and 30) and the exact parameter ranges vary slightly from one data set to another, depending on training time and specifics of each data set. For the full list of parameters and example values, see table 2. See the code for details, referenced at the end of this subsection. Each of the three experimental models are paired with each random configuration: original, era splitting, and directional era splitting. In this way all three models are trained and evaluated over the same set of parameter values. In each case, the model is trained on the the training set and evaluated on the OOS test set. Evaluation statistics are recorded for each configuration. Accuracy is the key statistic for classification problems. The regressor is converted to a binary classifier by rounding the predictions to the nearest integer (0 or 1). For regression problems the mean-squared error (MSE) and Pearson correlation (Corr) are observed.
All the code for reproducing the experiments presented in this paper is available in the following two links. The first contains the forked version of Scikit-Learn[38] which implements the era splitting models, and the second contains all the notebooks to recreate the figures and experiments from this research.
Parameter Example Range
Column Sample by Tree 0, 0.1, 0.3, 0.5, 0.7, 0.9, 1
L2 Regularization 0, 0.2, 0.4, 0.6, 0.8, 1
Learning Rate 0.01, 0.05, 0.1, 0.5, 1.0
Max Number of Bins 3, 4, 5, 7, 9
Max Depth 2, 3, 4, 5, 7, 9, 15
Max Number of Leaves 5, 7, 10, 16, 32
Min Child Samples 1, 3, 5, 10, 20
Boosting Rounds 5, 10, 20, 50, 100, 150
Split Type Original, Era Split, Dir. Era Split
Boltzmann Alpha -2, -1, 0, 1, 2
Table 2: Caption
5.1 The Shifted Sine Wave Data Set
The first experiment is a proof of concept for invariant learning in a simple regression setting. There is a 1-dimensional input feature which has an invariant aspect that is obscured by two types of randomness. One is a consistent Gaussian blur and the other is a random shift which is consistent inside each era but different across different eras of data. This part in particular simulates the distributional shift in the data generation process from one era to another.
The number of eras (environments) and the number of data points per era are pre-defined. The 1-dimensional input data is a random number on the interval $[0,2\pi]$. The target variable is a function of the the input. The invariant part of the target is a sine wave plus standard Gaussian noise. The environmental distribution shift is realized through a vertical shift of the target, where the magnitude of the shift is random, but consistent throughout each era. In this way, each era of data is a blurry sine wave that is shifted by a random amount. Pooling many eras of data together into one data set forms a cloud of data where the invariant mechanism, the sine wave, fades away and is not visually apparent. The experimental data set consists of 8 eras of data, each era having 64 data points.
Figure 4: A visual description of the data generation process for the shifted sine wave data set. Each era of training data starts with a sine wave (green), adds a random vertical shift and a random blur (Gaussian noise).
Evaluation metrics are computed against the test data, which is randomly generated by an additional era.
5.2 Synthetic Memorization Data Set
The synthetic memorization data set was developed in ([37]) specifically to investigate an ML model’s ability to learn invariant predictors in high-dimensional data in the presence of confounding spurious signals. It is a supervised binary classification task with target labels taking the values $\{0,1\}$. The data set is composed such that the first two input dimensions encode the invariant signal, which is an interlaced spiral shape, where one of the arms belongs to one class, and the other arm the other class. The remaining input dimensions are used to encode simple shortcuts, or spurious easy-to-learn signals. These are simply two clusters of data points in high dimensions, where one cluster belongs to one class and the other the other class. The clusters are linearly separable, being easy to learn. The spurious aspect is realized through a mechanism where each era (environment) of data has a shift in the location of the two clusters. The shift happens in such a way that, when all the environments of the training data set are pooled together, the shortcut is still present. Refer to figure 5 or figure 5 from ([37]) for a visualization of the data in four dimensions. When environments 1 & 2 are pooled together, there is still an easily identifiable linear decision boundary present in the spurious signal feature dimensions.
Figure 5: The synthetic memorization data set from ([37]). Each environment contains a random sample of the invariant spiral signal along the first two dimensions. The third and forth dimension contain a simple linearly-separable spurious signal that changes with each environment. When pooling environments together, the pooled spurious signals still create a simple linear decision boundary. During OOS testing, the spurious signal is replaced by random noise.
During test time the spurious signal is replaced by random noise. Any model that bases its predictions solely on the spurious signal will be worthless when evaluated on the test data. If a model has actually learned the invariant, albeit harder-to-learn signal, it will still perform well on the test set. The authors of ([37]) developed a technique called the AND-mask, which is an alteration to the naive gradient descent algorithm, incorporating the era-wise information. They were able to show that a naive MLP model simply learns the spurious signals in the data, and performs poorly on the test data. Meanwhile their AND-mask model was able to learn in such a way where it still performs well (98% accuracy) on the test data.
The training data contains 12,288 data points, each having 18 dimensions, coming from 16 eras. The test set is made of 2,000 data points. In order to use a regression model as a classifier, the model is trained on the classification labels $\{0,1\}$. Predictions are then rounded to the nearest integers, which is either a 00 or a $1$. The accuracy of the predictions can then be computed in the usual way according to 2-class classification problems.
5.3 Camelyon17 Data Set
The Camelyon17 data set ([23]) is an example of a real-world dataset exhibiting distributional shifts. This is one of the Wilds data sets that is distributed and supported by a team at Stanford ([33]). This data set aims at the automation of breast cancer detection in medical settings. The input data is a $96\times 96$ pixel histopathological image and the target label is a binary indicator that is positive if the central $32\times 32$ pixel region contains any tumor tissue. There are variations in data generation and processing technique from one hospital to another. The domain generalization problem is to learn to detect tumors from images from a number of hospitals, and have this detection ability transfer to new hospitals not in the training set.
Accompanying each input data point, there is an integer identifier corresponding to the hospital. This identifier is used as the training era (environment). The raw size of the image has $96\times 96\times 3=27,648$ features, which was taxing the memory capabilities of the research computer. To reduce the memory overhead the length and width of the inputs are reduced by a factor of 1/3 using the Wilds API, reducing the number of features to a more manageable $3,072$. There are $302,436$ data points in the training set coming from $3$ hospitals. The validation set contains $34,904$ data points from one hospital and the test set contains $85,054$ data points from one hospital. The Wilds data set provides established train, validation and test sets. Each model configuration is trained on the training set, and evaluated on the test set.
5.4 Numerai Data Set
Numerai is a San Francisco-based hedge fund which operates via a daily data science tournament in which the company distributes free, obfuscated data to participants, and those participants provide predictions back to Numerai. The predictions represent expected alpha, which is a directional up/down, bull/bear style prediction. On any day, a tournament participant will need to produce predictions for 5,000 to 10,000 unique stocks. The predictions are then scored against the true realizations that transpire in the markets, which then become the basis for future target labels.
Numerai distributes a large data set designed for supervised learning. Each row of data represents one stock at one time (era). The rows consist of many input feature columns, and also several target (output) columns. This research uses the main scoring target, called Cyrus. The Cyrus target values correspond to the optimal expected result over the subsequent 20 trading days, which is roughly equivalent to the stock’s future returns in the market subject to company-level risk requirements.
The raw time periods (eras) are numbered sequentially, starting at 1, and occur at weekly intervals. Each era represents one week of time. Currently there are over 1,060 eras of historical input and target pairs in the data set, and new eras are added each week, as the data becomes available. Each era contains more than 5,000 rows of data, on average, so the entire data set consists of over 5 million rows. The time horizon of the targets spans 4 weeks. Currently the full data set provides over 1,600 input features. In order to reduce training resources, a compact set of 244 features made popular by the tournament participant ShatteredX is used as our base feature set. It was shown with a baseline model that these features lead to out performance of the model which uses all the features.
For the experimental data, the entire data set is partitioned into 5 folds over time. The first 4 folds are used for training, the the fold identifier is used as the era identifier, in place of the true era. The last fold is used for evaluation. This process is important and effectively reduces the number of eras in the data from over 1,000 to 5. The terminology can get confusing, so these larger eras are called training eras. Each training eras spans about 4 years of time. This is a parameter choice that can reduce the complexity of the model. Reducing the number of training eras reduces the number of iterations in the algorithm, but also can lead to different performance capabilities. The current experiment did not explore fully the whole range of training era sizes, due to time constraints.
A notable grid search parameter is the number of estimators (boosting rounds), which varies between 50 and 1,000, to conserve training time. One ”benchmark” model well-known to the community is the LightGBM model with 2,000 trees, maximum depth of 5, 32 leaves, learning rate of 0.01 and column-sample-by-tree parameter of 0.1. This model is trained separately in order to make a comparison of our results to a well-known model.
The main evaluation metric is the era-wise correlation, which is called corr. The era-wise correlations are the mean Pearson correlations of the predictions with the targets computed over each era of data independently. A secondary important metric is called corr Sharpe, which is the mean era-wise correlation divided by the standard deviation of the era-wise scores.
6 Results
Experiment Original Era Split Dir. Era Split
Sine Wave (MSE) 1.79 1.71 1.75
Synth. Mem. (Acc) 52 88 96
Camelyon17 (Acc) 60 65 78
Numerai (Corr) 0.0236 0.0203 0.0265
Table 3: Best Test Scores by Model.
The results for all four experiments are summarized in table 3 and figure 6. Table 3 tabulates the best evaluation score over all the random configurations for each model split type. In each experiment the best model is either an era splitting model or a directional era splitting model. Figure 6 displays box plots of the evaluation metrics for both training and test sets over all the model configurations, separated by split type. A consistent trend emerges. The original split models are characterized by the best in-sample performance and the worst OOS performance, ie. the largest generalization gap. The era split and directional era split models have decreased in-sample and increased OOS performance, a smaller generalization gap, consistently in every experiment. This supports the claim of proposition 4.3, that these new splitting criteria act as effective regularizers.
Figure 6: Box plots of the evaluation metrics for each experiment, for training and test sets.
Figures 7 and 8 visualize the decision surfaces as a function of input data values for the shifted sine wave and synthetic memorization data sets. The sine wave experiment is one dimensional and easy to plot. The original split criterion results in a much noisier decision boundary than both era splitting and directional era splitting for this configuration. The synthetic memorization data set is high-dimensional, with the invariant swirl signal encoded on the first two dimensions. Plotting the predicted values as a function of these first two dimensions reveals whether or not the model has learned to use this signal. The original split model has been unable to learn the swirl signal, as well the predicted values are very extreme, meaning the model is very confident in those predictions. The high in-sample and low OOS accuracy scores indicate that these models have only learned the spurious signals which disappear in the OOS test set. On the other hand, both the experimental models have clearly learned the swirl signal, with higher predicted values along one arm and lower values along the other.
(a) Era Splitting
(b) Directional Era Splitting
Figure 7: Visualization of the decision boundary of the baseline model vs. the experimental model. On the left era splitting, on the right, directional era splitting. Both baseline and experimental model are trained with identical base parameters: boosting rounds = 100, max depth = 10, learning rate=1, and all other parameters default.
Both new splitting criteria improve OOS results on the Camelyon17 data set, indicating that these new criteria can be useful in real-world health related applications outside of finance.
The Numerai experiment shows that directional era splitting tends to consistently surpass the original splitting criterion, while the era splitting criterion did no surpass that of the original. In the direct comparison with the 2,000 tree benchmark model, the directional era splitting attained higher OOS corr with the target, 0.0271 compared too 0.0261 of the original. Results can be more nuanced than the bottom-line evaluation metric can show. In such a large parameter space, 15 to 20 configurations does not really explore a large portion of the space. To give more insight into the problem, there are a couple extra notebooks with experiments spanning more model parameters available in the online repository. In an additional experiment where the number of training eras is also varied from 3 to 8, several configurations of directional era splitting attained high correlation Sharpes, topping out at 1.26. That experiment can take 5 days or more to complete for 25 configurations.
All the results have been check-pointed in iPython notebooks available at the example code repository for this project:
github.com/jefferythewind/era-splitting-notebook-examples.
7 Discussion
The new splitting criteria presented in this research improve OOS performance and reduce in-sample performance of GBDT models, reducing the generalization gap in 4 concrete experiments. This supports the claimed roll as an effective regularizer. These experiments contain known and unknown distributional shifts within the training data and also between training and tests sets. Trees grown according to our new splitting criteria are more likely to generalize in these settings than trees grown according to the original splitting criterion.
Both of the new criteria are especially effective in the synthetic memorization problem and the Camelyon17 data set. The synthetic memorization contains a non-linear spiral signal as the invariant signal, while the true invariant signals present in the Camelyon17 data set are unknown. The positive results here showcase that there is a lot of information to lose if one does not take into account the domain (environmental) information.
The Numerai data set proved to be the most difficult to improve over the baseline, but the directional era splitting consistently did. It should be noted here that the era splitting criterion also tended to out-perform in preliminary experiments where a small grid was used, with a lower number of total trees. As the number of trees increased, the era splitting criterion itself appeared to become less effective. However there is still a lot left to explore here. While the era splitting correlation score ends up lower than the benchmark, other statistics can be much better, like correlation Sharpe and maximum draw down. The current Numerai experiment took 48 hours to train and evaluate 15 configurations. There are still many model configurations to experiment with. The number of training eras can also be varied and adapted, and this space is vast and relatively unexplored. An exploratory experiment showed some promise in this direction. In future experiments, the number of models will be expanded to several hundred, or even 1,000, which will take weeks or months to complete training. While out performance of individual models appears modest, the added variety from the the new split criteria can create the opportunity for more robust ensembles, which are averaged groups of models.
This research has taken a large step in the development of invariant learning for GBDTs, but is still in its infancy. Directional era splitting appears to work the best in real-world data sets, even though it strays farther from traditional ideas in decision trees than the era splitting criterion. Instead of variance-based impurity measures, this uses a directional-based measure, which is completely new to the field. This provides fresh ground for further theoretical interpretation and research.
Figure 8: Comparison of the decision boundaries between the experimental models on the first two dimensions of the synthetic memorization data set. The color of the data point shows the predicted value based on that input. The original model was unable to learn the swirl signal while era splitting and directional era splitting models were able to.
Time Complexity
One of the biggest obstacles to the wide-spread adoption of this model is the added time complexity that is added by the era splitting routine. Indeed, one of the main issues with GBDTs in general is that, in order to grow a tree, each node must compute the split score (equation 4) for every potential split in the entire data set. The calculation must be made for every distinct value for every feature in the data. The histogram method ([26]) in modern software libraries has helped by reducing the number of unique values in the features down to a set parameter value. Setting this parameter to a low value, like 5, can greatly reduce training time. The colsample bytree method also reduces the total number of computations needed by randomly sampling a subset of the features to use to build each tree, similar to the way bagging can reduce the number of rows.
In Era Splitting, the split gain must be computed not once but $M$ times per split, where $M$ is the number of training eras in our training data. For the Numerai data, $M$ can potentially take values up to 1,000. Run times for this implementation would take 1,000 times longer than the baseline model. This becomes prohibitive, as training times for larger models can easily take several days to complete training. In the experiment, the number of raw eras was reduced down to 5 training eras, which has a small time complexity and good performance. In other domains, such as the health domain characterizing the Camelyon17 data set, environments (such as hospitals) could be grouped together in logical ways, perhaps by geographic location. This, however, wasn’t an issue for Camelyon17 data, which only came from a handful of hospitals. More progress could be made to understand the trade offs, limitations, and performance affects pertaining to the number of training eras in the data.
Acknowledgments
I would like to express my gratitude to Richard Craib, Michael Oliver, and the entire Numerai research team for their invaluable assistance and support throughout this project. This project was funded in part through an internship with Numerai from the summer of 2022 to 2023.